OpenStudy (anonymous):

solve x-1 = sqrt(5x-9)

OpenStudy (mertsj):

Square both sides.

OpenStudy (anonymous):

x-1=sqrt(5x-9) (x-1)^2=5x-9 Then expand (x-1)^2 and post what you get

OpenStudy (anonymous):

x^2 - x - x + 1 ? @dfav

OpenStudy (anonymous):

yep yep

OpenStudy (anonymous):

not sure what to do with the x^2 now and how to get rid of it @dfav

OpenStudy (anonymous):

Here is what we have now x^2-x-x+1=5x-9 do you agree?

OpenStudy (anonymous):

yes that's right @dfav

OpenStudy (anonymous):

so it's be x=2 & x=5

OpenStudy (anonymous):

ok so combine all the like terms x^2-2x-5x+1+9=0

OpenStudy (anonymous):

x^2-7x+10=0

OpenStudy (anonymous):

which factors out to (x-5) (x-2) so...x=5 and x=2

OpenStudy (anonymous):

so which one do I know is the right answer, either 5 or 2 @dfav

OpenStudy (anonymous):

They are both possible solutions @Alexis2332 :)

OpenStudy (anonymous):

yup that's true- like I said earlier